What is the y-intercept of the graph of 3x-6y=7
Answer by algebrahouse.com(1659)

Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept

3x - 6y = 7
-6y = -3x + 7 {subtracted 3x form both sides}
y = (1/2)x - (7/6) {divided both sides by -6}
slope = 1/2
y-intercept = -7/6

For more help from me, visit: www.algebrahouse.com